OpenStudy (anonymous):

integrate e^x/sqrt(1-e^(2x))

OpenStudy (anonymous):

I thought it was using trig sub

OpenStudy (zale101):

You could use trig sub, that's actually more easier

OpenStudy (zale101):

Have you tried doing it?

OpenStudy (anonymous):

could u explain how to do it with trib sub

OpenStudy (anonymous):

it would tanx

OpenStudy (anonymous):

oh i get it

OpenStudy (anonymous):

thanks

OpenStudy (zale101):

Yes, tangent

OpenStudy (anonymous):

sqrt(sec ^2 -1)

OpenStudy (zale101):

U sub first u=e^x du=e^xdx \[\int\limits_{}^{}\frac{e^x}{1+e^{2x}}dx\] \[\int\limits_{}^{}\frac{du}{\sqrt{1+u^2}}\] Then Trig sub \(1+u^2=1+tan^2\theta=sec^2\theta\) \(tan\theta=u\) \(du=sec^2\theta ~d\theta\)

OpenStudy (anonymous):

thankk